2002 VIGRE REU PARTICIPANTS

Beitler, Lauren Beteta, Douglas Bucher, Russell (attending classes only) Butler, Emily Dhand, Vivek Dorn, Kimberly D’Souza, Keith Dudzik, Andrew Fitzpatrick, Andrew Liam Gelvin, Matthew Gostelow, Roli Hall, Diana Halper,Yehuda Holburt, Gabrielle Hou, Yuan Johnson, Paul Kalafarski, Adam Kilper, Gary Kribs, Robert Lee, Jungttwa Mercer, Joshua W. Morris, Colliness Nainani, Ajay Nicewicz, Maciej Ochsner, Evan Peters, Emily Pfaff, Catherine Purdy, David Purdy, Eric Redlich, Amanda E. Shaller, Alex G. Stover, Matthew Sullivan, George Yen-Chieh Tseng Tucker, Kevin F. Ward, Kenneth Welji, Shaffiq (reading only) Wheatcroft, David Whisenant, Houston L. White, Dane Wilson, Andrew Wolahan, Stephanie Wolff, Daniel M. Yin, Vera Zaraza, Derek